# Benchmarking Self-Supervised Methods for Accelerated MRI Reconstruction

### v1

@@ 2.1 Benchmarked methods
Table 1:
@@
MC - Eq. 3 dinv.loss.MCLoss()
@@ 2.1.2 Learning from multiple operators
A popular set of state-of-the-art approaches leverages the fact that the whole measurement space is spanned by multiple imaging operators. Measurement splitting methods, including SSDU (Yaman et al., 2020; Acar et al., 2021), randomly split the k-space into two sets at each iteration. One is used as the input and the other to construct the loss at the output:
@@
More recently, (Millard & Chiew, 2023) frames SSDU as a Bernoulli-noise special case of Noisier2Noise (Moran et al., 2019), and propose a modification to the SSDU loss that improves its results.
@@ 2.1.4 Data augmentation
EI is related to data augmentation (DA), which, when ground truth ${\mathbf{x}}_{\text{GT}}$ is available, can be used to constrain invariance or equivariance on ${\mathbf{x}}$. While (Fabian et al., 2021) applies this to MRI, VORTEX (Desai et al., 2022a) performs data augmentation on inputs ${\mathbf{y}}$ to attempt to solve the inverse problem.
@@ 2.1.5 Adversarial losses
${\mathcal{L}}_{\text{GAN}}=D(\tilde{\mathbf{A}}{f_{\theta}}({\mathbf{y}}),\tilde{\mathbf{y}})$ (8)
where $\tilde{\mathbf{y}}\sim\mathcal{Y}$ i.e. another randomly sampled measurement from the training dataset $\mathcal{Y}$. UAIR (Pajot et al., 2018) is a related method using the following loss:
@@ 2.1.6 Non-feedforward methods
Finally, we do not consider ground-truth-free methods that require inference or optimisation at inference-time, since these are impractical in real-world application, including Deep Image Prior methods (Darestani & Heckel, 2021), AmbientGAN (Bora et al., 2018), diffusion methods (Daras et al., 2023; Kawar et al., 2024), or implicit neural representation methods (Shen et al., 2024).
@@ 2.2.1 Noisy problems
In this paper, we separate the task of solving ill-posed problems with removing measurement noise. To achieve the latter, it is often possible to add a denoising term to the loss function. For example, Robust EI (Chen et al., 2022) adds the Stein’s Unbiased Risk Estimator (SURE) term. Noise2Recon (Desai et al., 2022b) and Robust SSDU (Millard & Chiew, 2024) both add a term similar to Noisier2Noise (Moran et al., 2019) to perform denoising on top of SSDU.
@@ 4 Experiments
Finally, we run further experiments to reproduce SSDU (Yaman et al., 2020) in Section 6.1 and test the robustness capabilities of our trained models to out-of-domain data in Section 6.2.
@@ 5 Results
Table 2: Test set results for 6x accelerated MRI on FastMRI brain dataset. Best unsupervised method in bold.
Figure 3: Sample reconstructions for 6x accelerated brain MRI reconstruction with average test-set PSNR labelled.
@@
As expected, measurement consistency (MC) cannot learn any information and recovers the no-learning performance. While EI and MOI both show good results compared to MC, our method, MO-EI, improves significantly on both these previous state-of-the-art methods, approaches the oracle supervised learning performance, and recovers details in the brain images where EI and MOI have strong artifacts (see insets). While SSDU qualitatively seems to remove some artifacts compared to no-learning, its metrics are very poor and drastically worse than no-learning; we explore why this is the case in Section 6.1.
@@ 6.1 Reproducing SSDU
We explore the seemingly poor performance of SSDU and attempt to reproduce the performance reported in (Yaman et al., 2020). First, we add 4 more unrolled iterations to the neural network from 3 to 7, increasing the training time by more-than-double. Note that this means these results are no longer directly comparable with other methods reported in Section 5. We observe a considerable qualitative performance increase but not quantitatively in Figs. 4 and 3. Then, noting that SSDU is trained on further subsampled measurements ${\mathbf{M}}_{1}{\mathbf{y}}_{i}$ but tested on just the subsampled measurements ${\mathbf{y}}_{i}$, we normalise the model output to the range $[0,1]$. This improves the results, but removes any quantitative information in the pixel values. Finally, we standardise the output to the mean and standard deviation of the ground truth ${\mathbf{x}}$. While this now shows an improvement over no-learning, this requires ground-truth statistics and thus would no longer be fully self-supervised. We note that the Noise2Inverse inference time procedure also does not help, differing from comparisons observed in dynamic MRI (Wang & Davies, 2024a).
@@
Table 3: Further results for training SSDU (Yaman et al., 2020) with a larger model (7 unrolled iters), and metrics after normalising output to $[0,1]$ or standardising output (requires ground truth statistics $\bar{\mathbf{x}}_{\text{GT}},\sigma_{\bar{\mathbf{x}}}$).
@@ Appendix B Data and hyperparameter details
For loss functions involving measurement splitting (i.e. SSDU), we use a 2D Gaussian splitting mask with $\rho=0.6$ following (Yaman et al., 2020). For losses involving equivariant imaging, we draw randomly a group transform at each iteration; for rotation the group being $G=\text{SO}(\mathbb{R}^{2})$ and for diffeomorphisms $G=\text{Diffeo}(\mathbb{R}^{2})$
@@ Appendix D Ablation
We ablate for the components of our proposed loss function by interpolating between existing methods MOI and EI and our method MO-EI, and report results in Table 5.

### v2

@@ 2.1 Benchmarked methods
Table 1: Self-supervised losses reviewed in Section 2.1, implemented in DeepInverse and benchmarked in our experiments.
@@
MC (Darestani & Heckel, 2021) Eq. 3 dinv.loss.MCLoss()
@@
Weighted-SSDU (Millard & Chiew, 2023) Eq. 7 dinv.loss.WeightedSplittingLoss()
Adversarial (Cole et al., 2021) Eq. 11 dinv.loss.UnsupAdversarialGeneratorLoss()
UAIR (Pajot et al., 2018) Eq. 12 dinv.loss.UAIRGeneratorLoss()
VORTEX (Desai et al., 2022a) Eq. 10 dinv.loss.VORTEXLoss()
EI (Chen et al., 2021) Eq. 9 dinv.loss.EILoss()
@@ 2.1.1 Measurement consistency
In the noisy scenario, some recent works such as ENSURE (Aggarwal et al., 2023) or GSURE (Metzler et al., 2020) generalise Stein’s Unbiased Risk Estimator (SURE) to reconstruct from noisy undersampled measurements. However, in the noiseless case considered here $\sigma\rightarrow 0$, these methods simply reduce to ${\mathcal{L}}_{\text{MC}}$, and so will not be compared.
@@ 2.1.2 Learning from multiple operators
A popular set of state-of-the-art approaches leverages the fact that the whole measurement space is spanned by multiple imaging operators. Measurement splitting methods, including SSDU and similar papers (Yaman et al., 2020; Acar et al., 2021; Huang et al., 2024; Zhou et al., 2022; Hu et al., 2021), randomly split the k-space into two sets at each iteration. One is used as the input and the other to construct the loss at the output:
@@
More recently, (Millard & Chiew, 2023) frames SSDU as a Bernoulli-noise special case of Noisier2Noise (Moran et al., 2019), and propose a weighting to the SSDU loss and an informed splitting strategy to improve SSDU:
${\mathcal{L}}_{\text{Weighted-SSDU}}$ $=(1-\mathbf{K})^{-1/2}{\mathcal{L}}_{\text{SSDU}}$ (6)
$\mathbf{K}$ $=(\mathbb{I}_{n}-\tilde{\mathbf{P}}\mathbf{P})^{-1}(\mathbb{I}_{n}-\mathbf{P})$ (7)
where $\mathbf{P}=\mathbb{E}[{\mathbf{M}}_{i}],\tilde{\mathbf{P}}=\mathbb{E}[{\mathbf{M}}_{1}]$ i.e. the PDFs of the imaging mask and the splitting mask, respectively.
@@ 2.1.4 Data augmentation
EI is related to data augmentation (DA), which, when ground truth ${\mathbf{x}}_{\text{GT}}$ is available, can be used to constrain invariance or equivariance on ${\mathbf{x}}$, applied to MRI in (Fabian et al., 2021). VORTEX (Desai et al., 2022a) builds on this in the semi-supervised context by performing data augmentation on the measurements when ground truth is not available:
${\mathcal{L}}_{\text{VORTEX}}=\begin{cases}{\mathcal{L}}_{\text{sup}}({f_{\theta}}({\mathbf{y}}),{\mathbf{x}}_{\text{GT}})&\text{if GT exists}\\ {\mathcal{L}}({\mathbf{T}}_{2}{f_{\theta}}(y,{\mathbf{A}}),{f_{\theta}}({\mathbf{T}}_{1}{\mathbf{y}},{\mathbf{A}}{\mathbf{T}}_{2}^{-1}))&\text{else}\end{cases}$ (10)
where ${\mathbf{T}}_{1}$ is a random k-space transform (e.g. add noise, phase shift) and ${\mathbf{T}}_{2}$ is a random image transform (e.g. shift, rotate). We consider the fully unsupervised case, replacing ${\mathcal{L}}_{\text{sup}}$ with ${\mathcal{L}}_{\text{MC}}$ to enforce measurement consistency, noting that we do not expect this to learn any new information as data in the null-space remains unobserved.
@@ 2.1.5 Adversarial losses
${\mathcal{L}}_{\text{adversarial}}=D(\tilde{\mathbf{A}}{f_{\theta}}({\mathbf{y}}),\tilde{\mathbf{y}})$ (11)
where $\tilde{\mathbf{y}}\sim\mathcal{Y}$ i.e. another randomly sampled ”real” measurement from the training dataset $\mathcal{Y}$. $D$’s aim is to distinguish between reconstructed measurements and training measurements, such that ${f_{\theta}}$’s aim is to reconstruct high quality measurements that are distinguishable from ”real” measurements. Note that in a non-generative context, it is unclear how this loss can learn the signal model and recover information from the null-space. UAIR (Pajot et al., 2018) is a related method using the following loss, originally proposed for inpainting:
@@ 2.1.6 Non-feedforward methods
Finally, we do not consider ground-truth-free methods that require lengthy iterative procedures or optimisation at inference-time, since these are impractical in real-world application, including Deep Image Prior methods (Darestani & Heckel, 2021), generative models e.g. AmbientGAN (Bora et al., 2018), diffusion methods (Daras et al., 2023; Kawar et al., 2024), implicit neural representation methods (Shen et al., 2024), or self-supervised plug-and-play training methods such as (Shafique et al., 2024).
@@ 2.2.1 Noisy problems
In this paper, we separate the task of solving ill-posed problems with removing measurement noise. To achieve the latter, it is often possible to add a denoising term to the loss function. For example, Robust EI (Chen et al., 2022) adds the Stein’s Unbiased Risk Estimator (SURE) term, which is adapted by ENSURE (Aggarwal et al., 2023). Noise2Recon (Desai et al., 2022b) and Robust SSDU (Millard & Chiew, 2024) both add a term similar to Noisier2Noise (Moran et al., 2019) to perform denoising on top of SSDU.
@@ 4 Experiments
Finally, we run further experiments to reproduce SSDU (Yaman et al., 2020) and its variants (Millard & Chiew, 2023) in Section 6.1 and test the robustness capabilities of our trained models to out-of-domain data in Section 6.2.
@@ 5 Results
SSDU* 18.27 .4876
Noise2Inverse* 18.49 .4693
Weighted-SSDU* 6.94 .2970
Adversarial 18.52 .4732
UAIR 14.00 .3715
VORTEX 27.75 .7898
@@
Table 2: Test set results for 6x accelerated MRI on FastMRI brain dataset. Best unsupervised method in bold. *we note the unexpected poor performance of these methods and further analyse them in Sections 6.1, 3 and 4.
Figure 3: Sample reconstructions for 6x accelerated brain MRI reconstruction with average test-set PSNR labelled, comparing a subset of methods from the full set of compared methods. See Table 2 for full results.
@@
As expected, measurement consistency (MC) cannot learn any information and recovers the no-learning performance. While EI and MOI both show good results compared to MC, our method, MO-EI, improves significantly on both these previous state-of-the-art methods, approaches the oracle supervised learning performance, and recovers details in the brain images where EI and MOI have strong artifacts (see insets).
While SSDU qualitatively seems to remove some artifacts compared to no-learning, its metrics are very poor and drastically worse than no-learning; we explore why this is the case in Section 6.1 and show how to improve the results. Additionally, the inference-time modification inspired by Noise2Inverse from Eq. 5 does not seem to improve the results. Similarly, the weighted version weighted-SSDU from Eq. 7 performs badly too, and we show how to retrieve better results in Section 6.1.
We note the very poor performance of the adversarial losses, and suggest that this is due to the very sensitive training of ${f_{\theta}}$ and $D$. Due to imbalanced training, the network seems to severely decimate the input far beyond the no-learning reconstruction, and we struggle to reproduce the results from (Cole et al., 2021), which may require heavy hyperparameter tuning and specialised model architectures; UAIR (Pajot et al., 2018) does not show results on accelerated MRI and we assume that their results do not transfer to this setting. Finally, VORTEX learns only marginal information compared to MC as expected, see Appendix B for a full analysis.
@@ 6.1 Reproducing SSDU
We explore the seemingly poor performance of SSDU and attempt to reproduce the performance reported in (Yaman et al., 2020). First, we add 4 more unrolled iterations to the neural network from 3 to 7, increasing the training computational cost by more-than-double. Note that this means these results are no longer directly comparable with other methods reported in Section 5. We observe a considerable qualitative performance increase but not quantitatively in Figs. 4 and 3. Then, noting that SSDU is trained on further subsampled measurements ${\mathbf{M}}_{1}{\mathbf{y}}_{i}$ but tested on just the subsampled measurements ${\mathbf{y}}_{i}$, we normalise the model output to the range $[0,1]$. This improves the results, but removes any quantitative information in the pixel values. Finally, we standardise the output to the mean and standard deviation of the ground truth ${\mathbf{x}}$. While this now shows an improvement over no-learning, this requires ground-truth statistics and thus would no longer be fully self-supervised. We note that the Noise2Inverse inference time procedure also does not help, differing from comparisons observed in dynamic MRI (Wang & Davies, 2024a).
To explore the performance of weighted-SSDU, we do similar and increase the number of unrolled iterations. This massively boosts the performance of weighted-SSDU and it quantitatively outperforms the no-learning reconstruction, suggesting that although the theory holds agnostic of the neural network architecture, in practice, its performance highly depends on it.
@@
Weighted-SSDU (3) 6.94 .2970
Weighted-SSDU (7) 29.44 .7822
Table 3: Further results for training SSDU (Yaman et al., 2020; Millard & Chiew, 2023) with a larger model (7 unrolled iters), and metrics after normalising output to $[0,1]$ or standardising output (requires ground truth statistics $\bar{\mathbf{x}}_{\text{GT}},\sigma_{\bar{\mathbf{x}}}$).
@@ 7 Conclusion
8 Acknowledgements
We thank Keying Guo for help with code implementation for reproducing (Millard & Chiew, 2023, 2024).
@@ Appendix B Data and hyperparameter details
For loss functions involving measurement splitting (i.e. SSDU), we use a 2D Gaussian splitting mask with $\rho=0.6$ following (Yaman et al., 2020), and for the weighted Noisier2Noise variant of SSDU we use an independent 1D Gaussian mask following (Millard & Chiew, 2023). For losses involving equivariant imaging, we draw randomly a group transform at each iteration; for rotation the group being $G=\text{SO}(\mathbb{R}^{2})$ and for diffeomorphisms $G=\text{Diffeo}(\mathbb{R}^{2})$. For VORTEX (Desai et al., 2022a), noting that the results of their various variants are very similar, for ${\mathbf{T}}_{1}$ we use random noise and their random phase errors with $\sigma=\alpha=1$ and for ${\mathbf{T}}_{2}$ we use random maximum ±$\pm 10\%$ shifts and random ±$\pm 15^{\circ}$ rotations. We also observe reduced performance when adding more transformations including full ±$\pm 360^{\circ}$ rotation, scaling, or shearing. We observe that when using the VORTEX consistency term on its own without any supervised or MC loss, it learns the trivial ${f_{\theta}}(y)=\mathbf{0}$, showing the superfluity of VORTEX.
For the adversarial losses, we use the same generator ${f_{\theta}}$ and the simple convolutional discriminator with skip connections used in (Cole et al., 2021).
@@ Appendix D Ablation
We ablate for the components of our proposed loss function by interpolating between existing methods MOI and EI and our method MO-EI, and report results in Table 5. We note that the base MOI and EI have similar results, and adding the two independent components of a) MO-EI and b) the diffeomorphic transforms helps, but combining these two provides the best performance in our proposed method, as per our theoretical framework.